# Transfer of quantum game strategies

### v1

@@ 6.4. An algebraic approach
(29) $\sum_{a\in A}p_{xx^{\prime},yy^{\prime}}^{aa,bb^{\prime}}=0,\;\;\;\;x\neq x^{\prime}\in X.$
Similarly,
(30) $\sum_{b\in B}p_{xx^{\prime},yy}^{aa^{\prime},bb}=\sum_{b\in B}p_{xx^{\prime},y^{\prime}y^{\prime}}^{aa^{\prime},bb},\;\;\;\;\text{ for \;}y,y^{\prime}\in Y$
@@
(31) $\sum_{b\in B}p_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb}=0,\;\;\;\;y\neq y^{\prime}\in Y.$
Finally, we require
(32) $\sum\limits_{a\in A,b\in B}p_{xx^{\prime},yy^{\prime}}^{aa,bb}=\delta_{xx^{\prime}}\delta_{yy^{\prime}}1,\;\;\;\;x,x^{\prime}\in X,y,y^{\prime}\in Y.$
Theorem 6.11.
Let $X_{i},A_{i},i=1,2$ be finite sets. Non-trivial *-representations of $\mathcal{C}_{X_{2},A_{2}}$ can be constructed from non-trivial *-representations of ${\rm C}^{*}$-algebras $\mathcal{C}_{X_{1},A_{1}}$ and $\mathcal{C}_{X_{2}A_{1},X_{1}A_{2}}$.
Proof.
Suppose there exists a Hilbert space $H$ and a non-trivial *-representation $\pi:\mathcal{C}_{X_{1},A_{1}}\rightarrow\mathcal{B}(H)$. Let $E=(\pi(e_{x_{1}x_{1}^{\prime},a_{1}a_{1}^{\prime}}))\in(M_{X_{1}A_{1}}\otimes\mathcal{B}(H))^{+}$ be the corresponding stochastic operator matrix (via [38, Theorem 5.2]). Similarly, suppose there exists a Hilbert space $K$ and non-trivial *-representation $\rho:\mathcal{C}_{X_{2}A_{1},X_{1}A_{2}}\rightarrow\mathcal{B}(K)$. Let $P:=(\rho(p_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1}^{\prime},a_{2}a_{2}^{\prime}}))\in(M_{X_{2}A_{1}X_{1}A_{2}}\otimes\mathcal{B}(K))^{+}$ be the corresponding operator matrix formed by applying $\rho$ to the generators of the universal ${\rm C}^{*}$-algebra. We claim that $P$ is a strongly stochastic operator matrix over $(X_{2}A_{1},X_{1}A_{2})$. Let
$P_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1}^{\prime},a_{2}a_{2}^{\prime}}:=\rho(p_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1}^{\prime},a_{2}a_{2}^{\prime}}),\;\;\;\;x_{i},x_{i}^{\prime}\in X_{i},a_{i},a_{i}^{\prime}\in A_{i},i=1,2.$
We see
${\rm Tr}_{X_{1}A_{2}}(P)$ $=$ $\sum\limits_{\begin{subarray}{c}x_{2},x_{2}^{\prime}\\ a_{1},a_{1}^{\prime}\end{subarray}}\sum\limits_{x_{1},a_{2}}\epsilon_{x_{2}x_{2}^{\prime}}\otimes\epsilon_{a_{1}a_{1}^{\prime}}\otimes P_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1},a_{2}a_{2}}$
$=$ $\sum\limits_{x_{2},x_{2}^{\prime}}\sum\limits_{a_{1},a_{1}^{\prime}}\epsilon_{x_{2}x_{2}^{\prime}}\otimes\epsilon_{a_{1}a_{1}^{\prime}}\otimes\delta_{x_{2}x_{2}^{\prime}}\delta_{a_{1}a_{1}^{\prime}}I_{K}$
$=$ $\sum\limits_{x_{2},a_{1}}\epsilon_{x_{2}x_{2}}\otimes\epsilon_{a_{1}a_{1}}\otimes I_{K}$
$=$ $I_{X_{2}A_{1}}\otimes I_{K}.$
Thus, $P$ is stochastic. Furthermore, if $\sigma_{X}=\sum_{x,x^{\prime}}\sigma_{xx^{\prime}}\epsilon_{xx^{\prime}}\in S(M_{X_{2}})$ is an arbitrary state, we see
${\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P))$ $=$ $\sum\limits_{\begin{subarray}{c}a_{1},a_{1}^{\prime}\\ a_{2},a_{2}^{\prime}\end{subarray}}\sum\limits_{\begin{subarray}{c}x_{2},x_{2}^{\prime}\\ x_{1}\end{subarray}}\sigma_{x_{2}^{\prime}x_{2}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes\epsilon_{a_{2}a_{2}^{\prime}}\otimes P_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1},a_{2}a_{2}^{\prime}}$
$=$ $\sum\limits_{\begin{subarray}{c}a_{1},a_{1}^{\prime}\\ a_{2},a_{2}^{\prime}\end{subarray}}\sum\limits_{x_{1},x_{2}}\sigma_{x_{2}x_{2}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes\epsilon_{a_{2}a_{2}^{\prime}}\otimes P_{x_{2}x_{2},a_{1}a_{1}^{\prime}}^{x_{1}x_{1},a_{2}a_{2}^{\prime}},$
where the latter equality follows from (28) and (29) in $\mathcal{C}_{X_{2}A_{1},X_{1}A_{2}}$. Let
$P_{a_{1}a_{1}^{\prime}}^{a_{2}a_{2}^{\prime}}:=\sum\limits_{x_{1}}P_{x_{2}x_{2},a_{1}a_{1}^{\prime}}^{x_{1}x_{1},a_{2}a_{2}^{\prime}},\;\;\;\;a_{i},a_{i}^{\prime}\in A_{i},i=1,2.$
We see
${\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P))$ $=$ $\sum\limits_{x_{2}}\sum\limits_{\begin{subarray}{c}a_{1},a_{1}^{\prime}\\ a_{2},a_{2}^{\prime}\end{subarray}}\sigma_{x_{2}x_{2}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes\epsilon_{a_{2}a_{2}^{\prime}}\otimes P_{a_{1}a_{1}^{\prime}}^{a_{2}a_{2}^{\prime}}$
$=$ $\sum\limits_{a_{1},a_{1}^{\prime}}\sum\limits_{a_{2},a_{2}^{\prime}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes\epsilon_{a_{2}a_{2}^{\prime}}\otimes P_{a_{1}a_{1}^{\prime}}^{a_{2}a_{2}^{\prime}}.$
Clearly, this shows
${\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P))={\rm Tr}_{X_{1}}(L_{\sigma_{X}^{\prime}}(P)),\;\;\;\;\sigma_{X},\sigma_{X}^{\prime}\in S(M_{X_{2}}).$
Furthermore, ${\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P))\in(M_{A_{1}A_{2}}\otimes\mathcal{B}(K))^{+}$ for any state $\sigma_{X}\in S(M_{X_{2}})$, as the slice of a positve element. To finish checking stochasticity, using (32) we see
${\rm Tr}_{A_{2}}({\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P)))$ $=$ $\sum\limits_{a_{2}}\sum\limits_{a_{1},a_{1}^{\prime}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes P_{a_{1}a_{1}^{\prime}}^{a_{2}a_{2}}$
$=$ $\sum\limits_{a_{1},a_{1}^{\prime}}\epsilon_{a_{1}a_{1}^{\prime}}\otimes\delta_{a_{1}a_{1}^{\prime}}I_{K}$
$=$ $\sum\limits_{a_{1}}\epsilon_{a_{1}a_{1}}\otimes I_{K}$
$=$ $I_{A_{1}}\otimes I_{K}.$
Thus, ${\rm Tr}_{X_{1}}(L_{\sigma_{X}}(P))$ is stochastic over $(A_{1},A_{2})$ for any $\sigma_{X}\in S(M_{X_{2}})$. By symmetry,
${\rm Tr}_{A_{2}}(L_{\sigma_{A}}(P))={\rm Tr}_{A_{2}}(L_{\sigma_{A}^{\prime}}(P)),\;\;\;\;\sigma_{A},\sigma_{A}^{\prime}\in S(M_{A_{1}}),$
where the latter is stochastic over $(X_{2},X_{1})$. This establishes our claim to strong stochasticity of $P$.
Define $F:=L_{J_{X_{1}}\otimes J_{A_{1}}}(E\otimes P)=(F_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}})$. Clearly, $F\in(M_{X_{2}A_{2}}\otimes\mathcal{B}(H\otimes K))^{+}$, and for $x_{2},x_{2}^{\prime}\in X_{2}$,
$\sum_{a_{2}}F_{x_{2}x_{2}^{\prime},a_{2}a_{2}}$ $=$ $\sum\limits_{a_{2}}\sum\limits_{\begin{subarray}{c}x_{1},x_{1}^{\prime}\\ a_{1},a_{1}^{\prime}\end{subarray}}E_{x_{1}x_{1}^{\prime},a_{1}a_{1}^{\prime}}\otimes P_{x_{2}x_{2}^{\prime},a_{1}a_{1}^{\prime}}^{x_{1}x_{1}^{\prime},a_{2}a_{2}}$
$=$ $\sum\limits_{x_{1},x_{1}^{\prime}}\sum\limits_{a_{1}}E_{x_{1}x_{1}^{\prime},a_{1}a_{1}}\otimes P_{x_{2}x_{2}^{\prime}}^{x_{1}x_{1}^{\prime}}$
$=$ $\sum\limits_{x_{1},x_{1}^{\prime}}\delta_{x_{1}x_{1}^{\prime}}I_{H}\otimes P_{x_{2}x_{2}^{\prime}}^{x_{1}x_{1}^{\prime}}$
$=$ $\sum\limits_{x_{1}}I_{H}\otimes P_{x_{2}x_{2}^{\prime}}^{x_{1}x_{1}}$
$=$ $I_{H}\otimes\delta_{x_{2}x_{2}^{\prime}}I_{K}.$
This implies ${\rm Tr}_{A_{2}}(F)=I_{X_{2}}\otimes I_{H}\otimes I_{K}$, showing that $F$ is stochastic over $(X_{2},A_{2})$. Using [38, Theorem 5.2] once more, this means we can ensure the existence of a *-representation $\tilde{\pi}:\mathcal{C}_{X_{2},A_{2}}\rightarrow\mathcal{B}(H\otimes K)$ such that $\tilde{\pi}(e_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}})=F_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}}$. Therefore, $\mathcal{C}_{X_{2},A_{2}}$ is non-trivial as a ${\rm C}^{*}$-algebra. ∎
Remark 6.12.
We may use a similar construction from the proof of Theorem 6.11 to show the existence of a surjective *-homomorphism $\pi:\mathcal{C}_{X_{2},A_{2}}\rightarrow\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{X_{1},A_{1}}$. Indeed: for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$ let
The next result is technical, showing that SQNS correlations have well-defined marginal QNS channels.
Proposition 6.13.
Every SQNS correlation $\Gamma:M_{X_{2}X_{2}\times A_{1}A_{1}}\rightarrow M_{X_{1}X_{1}\times A_{2}A_{2}}$ has well-defined left and right marginal QNS correlations $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}:M_{X_{2}X_{2}}\rightarrow M_{X_{1}X_{1}}$ and $\Gamma_{\rho}^{A_{1}A_{1}\rightarrow A_{2}A_{2}}:M_{A_{1}A_{1}}\rightarrow M_{A_{2}A_{2}}$ for any state $\sigma\in M_{A_{1}A_{1}},\rho\in M_{X_{2}X_{2}}$.
Proof.
For the moment, fix a state $\sigma\in M_{A_{1}A_{1}}$, and define the mapping $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}:M_{X_{2}X_{2}}\rightarrow M_{X_{1}X_{1}}$ via
We note that $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}$ is well-defined and completely positive, as $\Gamma$ is a well-defined quantum channel, $\sigma$ is a state, and taking the partial trace is a completely positive operation. Furthermore, as the partial trace is a trace-preserving operation, $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}$ is a quantum channel. By definition of an SQNS correlation, we automatically see that $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}$ is QNS. We now claim that $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}=\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\tilde{\sigma}}$ for all states $\sigma,\tilde{\sigma}\in M_{A_{1}A_{1}}$. Indeed: for $\rho\in M_{X_{2}X_{2}}$, as ${\rm Tr}_{A_{1}}(\sigma)={\rm Tr}_{A_{1}}(\tilde{\sigma})=1$, then ${\rm Tr}_{A_{1}}(\sigma-\tilde{\sigma})=0$. Using that $\Gamma$ is SQNS, we see
$\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}(\rho)-\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\tilde{\sigma}}(\rho)$ $=$ ${\rm Tr}_{A_{2}A_{2}}\bigg{(}\Gamma(\rho\otimes\sigma)-\Gamma(\rho\otimes\tilde{\sigma})\bigg{)}$
$=$ ${\rm Tr}_{A_{2}A_{2}}(\Gamma(\rho\otimes(\sigma-\tilde{\sigma}))=0.$
As $\rho\in M_{X_{2}X_{2}}$ was arbitrary, this shows $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}=\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\tilde{\sigma}}$, for any states $\sigma,\tilde{\sigma}\in M_{A_{2}A_{2}}$. A similar argument can be made showing the existence of well-defined left marginal channel $\Gamma_{\rho}^{A_{1}A_{1}\rightarrow A_{2}A_{2}}$ for corresponding state $\rho\in M_{X_{2}X_{2}}$. ∎
As (the proof of) Proposition 6.13 shows that the marginal QNS channels are the same no matter what choice of input state, we let $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}$ and $\Gamma^{A_{1}A_{1}\rightarrow A_{2}A_{2}}$ denote the right (respectively, left) marginal channel of $\Gamma$, for an arbitrary fixed state $\sigma\in M_{A_{1}A_{1}}$ (respectively, $\rho\in M_{X_{2}X_{2}}$).
Remark 6.14.
Note that, in the event $\Gamma\in\mathcal{C}_{\rm st}$ for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$, then the marginal channels obtained from Proposition 6.13 (when considering $\Gamma$ inside $\mathcal{Q}_{\rm st}$) coincide with the marginal channels discussed in [19, Section 5].

### v2

@@ 6.1. Representations of SQNS correlations via operator systems
Thus,
by the injectivity of the minimal operator system tensor product, $C\in(\mathcal{L}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{L}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{L}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{L}_{B_{1},B_{2}})^{+}$. By [17, Proposition 1.9], and [39, Proposition 5.5],
Arguing now as in [39, Theorem 6.2], this establishes the claim.
(ii) First suppose that
Applying the GNS construction to $\tilde{s}$ and using [39, Theorem 5.2], we obtain a Hilbert space $H$, a unit vector $\xi\in H$ and mutually commuting stochastic operator matrices
satisfying
acting on $K$ so that $\Gamma=\Gamma_{E,F,\eta}$, for $E=E_{X}\cdot E_{Y}$ and $F=F_{A}\cdot F_{B}$. Let $\pi_{X},\pi_{Y},\pi_{A}$ and $\pi_{B}$ be the (unital) *-representations of $\mathcal{C}_{X_{2},X_{1}},\mathcal{C}_{Y_{2},Y_{1}},\mathcal{C}_{A_{1},A_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$ on $\mathcal{B}(K)$ arising from $E_{X},E_{Y},F_{A}$, and $F_{B}$ respectively (see [39, Theorem 5.2]). Then $\pi:=\pi_{X}\otimes\pi_{Y}\otimes\pi_{A}\otimes\pi_{B}$ is a unital *-representation of $\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{Y_{2},Y_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{B_{1},B_{2}}$ on $\mathcal{B}(K)$. Using Theorem 6.4 once more, if we let $s$ be the restriction to $\mathcal{T}_{X_{2},X_{1}}\otimes_{\rm c}\mathcal{T}_{Y_{2},Y_{1}}\otimes_{\rm c}\mathcal{T}_{A_{1},A_{2}}\otimes_{\rm c}\mathcal{T}_{B_{1},B_{2}}$ of the state
$w\mapsto\langle\pi(w)\eta,\eta\rangle,$
it is clear that $\Gamma=\Gamma_{s}$.
$w\mapsto\langle\pi(w)\tilde{\eta},\tilde{\eta}\rangle$,
Now let $s:\mathcal{T}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{T}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{T}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{T}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ be a state, and using the fact that the minimal operator system tensor product is injective (see [20]) let $\tilde{s}:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{C}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ be an extension of $s$. By [24, Corollary 4.3.10], $\tilde{s}$ can be approximated in the $\rm weak$$^{*}$ topology by elements of the convex hull of vector states on $\pi_{X}(\mathcal{C}_{X_{2},X_{1}})\otimes_{\rm min}\pi_{Y}(\mathcal{C}_{Y_{2},Y_{1}})\otimes_{\rm min}\pi_{A}(\mathcal{C}_{A_{1},A_{2}})\otimes_{\rm min}\pi_{B}(\mathcal{C}_{B_{1},B_{2}})$ (where $\pi_{X},\pi_{Y},\pi_{A},$ and $\pi_{B}$ are unital *-representations of $\mathcal{C}_{X_{2},X_{1}},\mathcal{C}_{Y_{2},Y_{1}},\mathcal{C}_{A_{1},A_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$, respectively). Using an argument similar to the proof of [7, Theorem 5.6] or [39, Theorem 6.5], we can show that $\Gamma$ is a limit of quantum SQNS correlations.
(iv) This proof is along the lines of the proof for [39, Theorem 6.7]; we include the details for the convenience of the reader. We first let $s$ be a state on
By [26, Theorem 9.9] and [17, Proposition 1.9], we may consider $s$ as an element of
$\hskip-18.49411pt\bigg({\rm OMAX}(\mathcal{T}_{X_{2},X_{1}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{Y_{2},Y_{1}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{A_{1},A_{2}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{B_{1},B_{2}})\bigg)^{+}.$
$\Gamma_{s}=\sum\limits_{j=1}^{k}\lambda_{j}\Phi_{X}^{(j)}\otimes\Phi_{Y}^{(j)}\otimes\Phi_{A}^{(j)}\otimes\Phi_{B}^{(j)}.$ (27)
By Remark 3.11, this shows $\Gamma_{s}\in\mathcal{Q}_{\rm sloc}$.
Now suppose $\Gamma$ is of the form (27), with $s$ a functional on $\mathcal{T}_{X_{2},X_{1}}\otimes\mathcal{T}_{Y_{2},Y_{1}}\otimes\mathcal{T}_{A_{1},A_{2}}\otimes\mathcal{T}_{B_{1},B_{2}}$ such that $\Gamma=\Gamma_{s}$. Let $E_{X}^{(j)}\in(M_{X_{2}}\otimes M_{X_{1}})^{+}$ (resp. $E_{Y}^{(j)}\in(M_{Y_{2}}\otimes M_{Y_{1}})^{+},E_{A}^{(j)}\in(M_{A_{1}}\otimes M_{A_{2}})^{+}$, and $E_{B}^{(j)}\in(M_{B_{1}}\otimes M_{B_{2}})^{+}$) be the Choi matrix of $\Phi_{X}^{(j)}$ (resp. $\Phi_{Y}^{(j)},\Phi_{A}^{(j)}$ and $\Phi_{B}^{(j)}$); then $E_{X}^{(j)},E_{Y}^{(j)},E_{A}^{(j)}$, and $E_{B}^{(j)}$ are stochastic operator matrices acting on $\mathbb{C}$. By [39, Theorem 5.2], there exist positive functionals $\phi_{X}^{(j)}:\mathcal{T}_{X_{2},X_{1}}\rightarrow\mathbb{C},\phi_{Y}^{(j)}:\mathcal{T}_{Y_{2},Y_{1}}\rightarrow\mathbb{C},\phi_{A}^{(j)}:\mathcal{T}_{A_{1},A_{2}}\rightarrow\mathbb{C}$, and $\phi_{B}^{(j)}:\mathcal{T}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ such that
and thus, by Lemma 6.5 is a state on
as claimed. ∎
Remark 6.7.
We note that, using an almost identical argument as in (iv), we have an affine isomorphism between the state space of ${\rm OMIN}(\mathcal{S}_{X_{2},X_{1}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{Y_{2},Y_{1}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{A_{1},A_{2}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{B_{1},B_{2}})$ and $\mathcal{C}_{\rm sloc}$; the classical analogue for cases (i)-(iii) are addressed in [20, Theorem 7.11].
Corollary 6.8
The set $\mathcal{Q}_{\rm sqc}$ is closed and convex.
Proof.
[…]
Remark 6.9.
By [20, Theorem 7.11], we may use an identical argument as in Corollary 6.8 to show that $\mathcal{C}_{\rm sqc}$ is closed and convex.
Remark 6.10.
Theorem 6.11
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
Statement (i) follows from Remark 6.10, and an application of [39, Theorem 8.3]. Statement (ii) follows from (i), and Theorem $6.$ Statements (iii) and (iv) follow similarly, when paired with [39, Remark 8.1]. ∎
As in the setup before Theorem 6.6, for a linear functional
Theorem 6.12
The map $s\mapsto\Gamma_{s}$ is an affine isomorphism from
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
(i) Let $\Gamma\in\mathcal{Q}_{\rm sns}^{\rm bi}$. If
is the Choi matrix of $\Gamma$ and
for all $x_{i},x_{i}^{\prime}\in X,i=1,2$. Note that
$L_{\rho}(C)\in\mathcal{L}_{X}\text{ for all }\rho\in M_{YAB}.$
$L_{\rho}(C)\in\mathcal{L}_{Y}\text{ for all }\rho\in M_{XAB},$
$L_{\rho}(C)\in\mathcal{L}_{A}\text{ for all }\rho\in M_{XYB},$
@@
$L_{\rho}(C)\in\mathcal{L}_{B}\text{ for all }\rho\in M_{XYA}.$
(ii)-(iv) The proofs for (ii)-(iv) follow exactly the same as in the proofs of Theorem 6.6 (ii)-(iv), just by replacing the use of [39, Theorem 5.2] with [7, Theorem 3.4], the use of Remark 3.11 with Remark 3.13, and by replacing $\mathcal{T}_{X_{2},X_{1}}$ and $\mathcal{C}_{X_{2},X_{1}}$ (resp. $\mathcal{T}_{Y_{2},Y_{1}}$ and $\mathcal{C}_{Y_{2},Y_{1}}$, $\mathcal{T}_{A_{1},A_{2}}$ and $\mathcal{C}_{A_{1},A_{2}}$, $\mathcal{T}_{B_{1},B_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$) with $\mathcal{T}_{X}$ and $\mathcal{C}_{X}$ (resp. $\mathcal{T}_{Y}$ and $\mathcal{C}_{Y}$, $\mathcal{T}_{A}$ and $\mathcal{C}_{A}$, $\mathcal{T}_{B}$ and $\mathcal{C}_{B}$). ∎
We end this subsection with the following (obvious) result.
Corollary 6.13
The set $\mathcal{Q}_{\rm sqc}^{\rm bi}$ is closed and convex.
6.2. Concurrent strategies
The goal of this subsection is to apply the results connecting subclasses of strongly QNS correlations with state spaces on tensor products of canonical operator systems previously developed, to the study of a particular class of quantum input-output game which are called concurrent games. A quantum non-local game $\varphi:\mathcal{P}_{XX}\rightarrow\mathcal{P}_{AA}$ is concurrent if $\varphi(\tilde{J}_{X})=\tilde{J}_{A}$. A QNS correlation $\Phi\in\mathcal{Q}_{\rm ns}$ is called concurrent if $\Phi(\tilde{J}_{X})=\tilde{J}_{A}$; equivalently, if $\Phi$ is a perfect strategy for the (trivial) implication game $\varphi_{\tilde{J}_{X}\rightarrow\tilde{J}_{A}}$.